# Beyond the Hodge Theorem: curl and asymmetric pseudodifferential projections

### v3

@@ 2.1 Operator theoretic definition and main properties
In order to prove the above theorem, let us introduce an auxiliary elliptic operator: extended curl.
2.2 Extended curl
Definition 2.3.
We define extended curl to be the operator
$\operatorname{curl}_{E}:=\begin{pmatrix}\operatorname{curl}&\mathrm{d}\\ \delta&0\end{pmatrix}:(\Omega^{1}\cap H^{1})\oplus(\Omega^{0}\cap H^{1})\to\Omega^{1}\oplus\Omega^{0}\,.$ (2.6)
The operator (2.6) appeared in [42], see also [2, p. 229], [3, pp. 44–45 and p. 66], [4, p. 405], [50, pp. 190–191] and [22, §4].
Remark 2.4.
Note that $\operatorname{curl}_{E}$ is an operator of Dirac type, namely, its square is $\delta\mathrm{d}+\mathrm{d}\delta$ on $\Omega^{1}\oplus\Omega^{0}$ (minus the nonpostive Hodge Laplacian). Furthermore, (2.6) is the restriction of the signature operator $\mathrm{d}+\delta$ to differential forms of even degree, in the sense that
$(\operatorname{curl}_{E})_{\mathrm{prin}}(x,\xi)=\begin{pmatrix}[(\operatorname{curl})_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi)&i\xi_{\alpha}\\ -ig^{\beta\gamma}(x)\,\xi_{\gamma}&0\end{pmatrix}.$ (2.7)
The eigenvalues of (2.7) are
±$\pm\|\xi\|\,,$ (2.8)
each with multiplicity two, so that we have
$\det\left[(\operatorname{curl}_{E})_{\mathrm{prin}}\right]=\|\xi\|^{4}\,,$ (2.9)
Compare formulae (2.8) and (2.9) with (2.3) and (2.2) respectively.
Formulae (2.9) and (2.6) imply that $\operatorname{curl}_{E}$ is elliptic and self-adjoint.
Definition 2.5 ([47, Definition 1.7]).
$L(\mathcal{D}(L)\cap V)\subseteq V.$
Proposition 2.6.
$\Omega^{1}\oplus\Omega^{0}=[\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1}]\oplus[\delta\Omega^{2}\oplus 0]\oplus[\mathcal{H}^{1}\oplus\mathcal{H}^{0}]$ (2.10)
which are invariant subspaces of the operator $\operatorname{curl}_{E}$ in the sense of Definition 2.5. Accordingly, the operator $\operatorname{curl}_{E}$ decomposes into a direct sum of three operators
$\operatorname{curl}_{E}=\operatorname{curl}_{E,\mathrm{d}}\oplus\operatorname{curl}_{E,\delta}\oplus\operatorname{curl}_{E,\mathcal{H}}\,,$ (2.11)
where
$\operatorname{curl}_{E,\mathrm{d}}:(\mathrm{d}\Omega^{0}\cap H^{1})\oplus(\delta\Omega^{1}\cap H^{1})\to\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1},$ (2.12)
$\operatorname{curl}_{E,\delta}:(\delta\Omega^{2}\cap H^{1})\oplus 0\to\delta\Omega^{2}\oplus 0,$ (2.13)
$\operatorname{curl}_{E,\mathcal{H}}:\mathcal{H}^{1}\oplus\mathcal{H}^{0}\to\mathcal{H}^{1}\oplus\mathcal{H}^{0}.$ (2.14)
Proof.
That the three subspaces in the RHS of (2.10) are mutually orthogonal and their direct sum gives the whole space $\Omega^{1}\oplus\Omega^{0}$ follows immediately from the Hodge decomposition theorem. Let us show that they are invariant subspaces of $\operatorname{curl}_{E}\,$. We will do this in several steps.
Step 2. The facts that
  • $[\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1}]\oplus[\delta\Omega^{2}\oplus 0]$ is the orthogonal complement of $\mathcal{H}^{1}\oplus\mathcal{H}^{0}$
imply that $[\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1}]\oplus[\delta\Omega^{2}\oplus 0]$ is an invariant subspace of $\operatorname{curl}_{E}\,$.
$\lambda\in\mathbb{R}.$ (2.15)
$\lambda\neq 0.$ (2.16)
$0=\mu_{0}<\mu_{1}\leq\mu_{2}\leq\dots$ (2.17)
be the eigenvalues of $\,-\Delta\,$ enumerated in increasing order with account of multiplicities. It will be convenient for us to further restrict the choice of $\lambda$ by imposing the condition
$\lambda\not\in\{\pm\sqrt{\mu_{1}}\,,\,\pm\sqrt{\mu_{2}}\,,\dots\}.$ (2.18)
Step 5. In view of (2.15) the operator $\mathcal{R}(\lambda)$ is self-adjoint. The facts that
  • $[\mathrm{d}\Omega^{0}\oplus\delta\Omega^{1}]\oplus[\delta\Omega^{2}\oplus 0]$ is an invariant subspace of $\mathcal{R}(\lambda)$ and
$\mathrm{d}\Omega^{k-1}_{\infty}:=\{\mathrm{d}\omega\ |\ \omega\in\Omega^{k-1}_{\infty}\},$ (2.19)
$\delta\Omega^{k+1}_{\infty}:=\{\delta\omega\ |\ \omega\in\Omega^{k+1}_{\infty}\}.$ (2.20)
The Hilbert spaces $\Omega^{k}$, $\mathrm{d}\Omega^{k-1}$ and $\delta\Omega^{k+1}$ are defined as $L^{2}$ closures of the vector spaces $\Omega^{k}_{\infty}$, $\mathrm{d}\Omega^{k-1}_{\infty}$ and $\delta\Omega^{k+1}_{\infty}$ respectively, see [36, Corollary 3.4.2]. This implies that the vector space $\delta\Omega^{2}_{\infty}\oplus 0$ is dense in the Hilbert spaces $\delta\Omega^{2}\oplus 0$. As $\mathcal{R}(\lambda)$ is a bounded operator, in order to prove that $\delta\Omega^{2}\oplus 0$ is an invariant subspace of $\mathcal{R}(\lambda)$ it is sufficient to prove that
$[\mathcal{R}(\lambda)](\delta\Omega^{2}_{\infty}\oplus 0)\subseteq\delta\Omega^{2}_{\infty}\oplus 0\,.$ (2.21)
$v:=\mathcal{R}(\lambda)z=\begin{pmatrix}v_{1}\\ v_{0}\end{pmatrix}\in[\mathrm{d}\Omega^{0}_{\infty}\oplus\delta\Omega^{1}_{\infty}]\oplus[\delta\Omega^{2}_{\infty}\oplus 0].$ (2.22)
Here $z_{1}$, $v_{1}$ are 1-forms and $v_{0}$ is a 0-form (scalar). The $v_{1}$ and $v_{0}$ are infinitely smooth because the operator $\operatorname{curl}_{E}$ is elliptic. Formula (2.22) can be rewritten more explicitly as
$\operatorname{curl}v_{1}+\mathrm{d}v_{0}-\lambda v_{1}$ $=z_{1}\,,$ (2.23)
$\delta v_{1}-\lambda v_{0}$ $=0\,.$ (2.24)
We have [53, Section 6.8, p. 223]
$\Omega^{k}_{\infty}=\mathrm{d}\Omega^{k-1}_{\infty}\oplus\delta\Omega^{k+1}_{\infty}\oplus\mathcal{H}^{k},$ (2.25)
compare with (1.1). We already know (see Step 2) that the 1-form $v_{1}$ does not contain a contribution from harmonic 1-forms, so formulae (2.25), (2.19) and (2.20) give us
$v_{1}=v_{1\mathrm{d}}+v_{1\delta}\,,$ (2.26)
where
$v_{1\mathrm{d}}=\mathrm{d}w_{0}\,,$ (2.27)
$v_{1\delta}=\delta w_{2}$ (2.28)
Step 8. Applying the operator $\delta$ to (2.23) and using formulae (2.26)–(2.28) as well as the fact that $\delta z_{1}=0$, we get $\Delta(\lambda w_{0}-v_{0})=0$, which implies
$v_{0}=\lambda w_{0}+C,$ (2.29)
where $C\in\mathbb{R}$ is a constant. Substituting (2.26) and (2.29) into (2.24) and using formulae (2.27), (2.28), we get
$-\Delta w_{0}-\lambda^{2}w_{0}=\lambda C.$ (2.30)
In view of (2.16) and (2.18) equation (2.30) has the unique solution $w_{0}=-\lambda^{-1}C$, which immediately implies
$v_{0}=0,\qquad v_{1\mathrm{d}}=0.$ (2.31)
It only remains to observe that formulae (2.22), (2.26), (2.28) and (2.31) imply $v\in\delta\Omega^{2}_{\infty}\oplus 0$, thus proving the inclusion (2.21). ∎
2.3 Decomposition of the spectrum of extended curl
Recall that extended curl (2.6) is a self-adjoint elliptic operator, hence its spectrum is discrete and eigenfunctions infinitely smooth.
Proposition 2.7.
The spectrum of $\operatorname{curl}_{E}$ decomposes into three parts as
$\sigma(\operatorname{curl}_{E})=\sigma(\operatorname{curl})\cup\sigma(\operatorname{curl}_{E,\mathrm{d}})\cup\sigma(\operatorname{curl}_{E,\mathcal{H}}),$ (2.32)
where $\operatorname{curl}$, $\operatorname{curl}_{E,\mathrm{d}}$ and $\operatorname{curl}_{E,\mathcal{H}}$ are the operators (2.4), (2.12) and (2.14) respectively. Formula (2.32) is understood as the (disjoint) union of spectra with account of multiplicities.
Proof.
Proposition 2.7 is an immediate consequence of Proposition 2.6 and the observation that the operator $\,\operatorname{curl}_{E,\delta}\,$ defined by formulae (2.6) and (2.13) is precisely our original operator $\operatorname{curl}$ defined by formula (2.4). ∎
Proposition 2.7 tells us that the spectrum of curl sits inside the spectrum of extended curl and the two spectra differ by $\sigma(\operatorname{curl}_{E,\mathrm{d}})\cup\sigma(\operatorname{curl}_{E,\mathcal{H}})$. So the task at hand is working out the spectra of $\operatorname{curl}_{E,\mathrm{d}}$ and $\operatorname{curl}_{E,\mathcal{H}}\,$. The following two lemmata address this issue.
Lemma 2.8.
Proof.
Lemma 2.8 is an immediate consequence of formulae (2.6) and (2.14). ∎
Lemma 2.9.
We have
$\sigma(\operatorname{curl}_{E,\mathrm{d}})=\{\pm\sqrt{\mu_{1}}\,,\,\pm\sqrt{\mu_{2}}\,,\dots\},$ (2.33)
where the $\mu_{j}$ are the eigenvalues (2.17) of $\,-\Delta\,$.
Proof.
$\mathrm{d}v_{0}$ $=\lambda\mathrm{d}w_{0}\,,$ (2.34)
$-\Delta w_{0}$ $=\lambda v_{0}\,.$ (2.35)
Elementary analysis of (2.34) (2.35) yields (2.33). ∎
Proposition 2.7 and Lemmata 2.82.9 tell us that the spectrum of $\operatorname{curl}$ can be recovered from the spectrum of $\operatorname{curl}_{E}\,$, provided one knows the spectrum of the Laplace–Beltrami operator $\Delta$. Moreover, the eta functions of $\operatorname{curl}$ and $\operatorname{curl}_{E}\,$ are the same, because contributions to $\sigma(\operatorname{curl}_{E})$ coming from the Laplace–Beltrami operator are symmetric about zero.
2.4 Proof of Theorem 2.1
Proof of Theorem 2.1.
(a) Standard elliptic theory tells us that the operator $\operatorname{curl}_{E}$ defined by formula (2.6) is self-adjoint in the operator theoretic sense. Proposition 2.6 and formulae (2.6) and (2.13) tell us that our original operator $\operatorname{curl}$ defined by formula (2.4) is part of the operator $\operatorname{curl}_{E}\,$; here the concept of ‘part of an operator’ is understood in accordance with [47, Definition 1.7]. Self-adjointness of $\operatorname{curl}$ now follows from [47, Proposition 3.11].
(b) Discreteness of the spectrum of $\operatorname{curl}$ follows from Proposition 2.7. As to the statement that the spectrum of $\operatorname{curl}$ is unbounded both from above and from below, we prove it by writing down spectral asymptotics.
Let $\lambda_{k}$, $k\in\mathbb{Z}\setminus\{0\}$, be the eigenvalues of $\operatorname{curl}$. The choice of particular enumeration is irrelevant for our purposes, but what is important is that eigenvalues are enumerated with account of their multiplicity. We define the positive ($+$) and negative (--) counting functions for $\operatorname{curl}$ as
$N^{\pm}_{\operatorname{curl}}(\lambda):=\begin{cases}0\quad&\text{for}\quad\lambda\leq 0,\\ \sum_{0<\pm\lambda_{k}<\lambda}\,1\quad&\text{for}\quad\lambda>0.\end{cases}$ (2.36)
We define, in a similar fashion, counting functions $N^{\pm}_{\operatorname{curl}_{E}}(\lambda)$ and $N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)$ for the operators $\operatorname{curl}_{E}$ and $\operatorname{curl}_{E,d}$ respectively. Proposition 2.7 and Lemma 2.8 tell us that
$N^{\pm}_{\operatorname{curl}}(\lambda)=N^{\pm}_{\operatorname{curl}_{E}}(\lambda)-N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda).$ (2.37)
Lemma 2.9 tells us that the evaluation of $N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)$ reduces to the evaluation of eigenvalues of the Laplace–Beltrami operator. The latter is a subject which has been extensively studied over the last 100 years. Application of [33, Theorem 1.1] gives the following asymptotic formula for $N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)$ with sharp remainder term estimate:
$N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)=\frac{\operatorname{Vol}(M)}{6\pi^{2}}\lambda^{3}+O(\lambda^{2})\qquad\text{as}\qquad\lambda\to+\infty,$ (2.38)
where $\operatorname{Vol}(M)$ is the Riemannian volume of the manifold.
Application of [34, Theorem 0.1] gives the following asymptotic formula for $N^{\pm}_{\operatorname{curl}_{E}}(\lambda)$:
$N^{\pm}_{\operatorname{curl}_{E}}(\lambda)=\frac{\operatorname{Vol}(M)}{3\pi^{2}}\lambda^{3}+O(\lambda^{2})=2N^{\pm}_{\operatorname{curl}_{E,d}}(\lambda)+O(\lambda^{2})\qquad\text{as}\qquad\lambda\to+\infty.$ (2.39)
The factor 2 appears in the RHS of (2.39) because the eigenvalues of the principal symbol of the operator $\operatorname{curl}_{E}\,$, given by formula (2.8), have multiplicity two.
Formulae (2.37)–(2.39) imply Bär’s [7, Theorem 3.6] asymptotic formula
$N^{\pm}_{\operatorname{curl}}(\lambda)=\frac{\operatorname{Vol}(M)}{6\pi^{2}}\lambda^{3}+O(\lambda^{2})\qquad\text{as}\qquad\lambda\to+\infty.$ (2.40)
Formula (2.40) shows that the spectrum of $\operatorname{curl}$ accumulates to +$\infty$+ ∞ and $\infty$- ∞.
(c) Suppose that zero is an eigenvalue of $\operatorname{curl}$. Let
$u\in\delta\Omega^{2}_{\infty}$ (2.41)
$\mathrm{d}u=0.$ (2.42)
But (2.41) implies
$\delta u=0.$ (2.43)
Formulae (2.42) and (2.43) tell us that the 1-form $u$ is harmonic. As $\delta\Omega^{2}\cap\mathcal{H}^{1}=\{0\}$, we conclude that $u=0$, which is a contradiction.
(d) Proposition 2.7, Lemmata 2.8 and 2.9 and part (c) of Theorem 2.1 tell us that zero is an eigenvalue of $\operatorname{curl}_{E}$ of multiplicity $\operatorname{dim}\mathcal{H}^{1}+1$. Let $P_{\mathcal{H}^{1}}$ and $P_{\mathcal{H}^{0}}$ be orthogonal projections onto the spaces of harmonic 1-forms and 0-forms respectively. Put
@@ 3 Pseudodifferential operators acting on 1-forms
$q_{\alpha}{}^{\beta}(x,\xi)\sim[q_{s}]_{\alpha}{}^{\beta}(x,\xi)+[q_{s-1}]_{\alpha}{}^{\beta}(x,\xi)+\dots$ (3.2)
be its representation in local coordinates (the same for $x$ and $y$). Here $\sim$ stands for asymptotic expansion [49, § 3.3], and ‘polyhomogeneous’ means that $[q_{s-k}]_{\alpha}{}^{\beta}(x,\lambda\,\xi)=\lambda^{s-k}\,[q_{s-k}]_{\alpha}{}^{\beta}(x,\xi)$ for any positive $\lambda$ and $k=0,1,2,\dots$.
@@
$[Q_{\mathrm{prin}}]_{\alpha}{}^{\beta}(x,\xi):=[q_{s}]_{\alpha}{}^{\beta}(x,\xi)$ (3.3)
@@
However, the subleading term, $[q_{s-1}]_{\alpha}{}^{\beta}\,$, is not invariant under change of coordinates. The task at hand is to define and provide a formula for the subprincipal symbol of $Q$, by determining appropriate correction terms to $[q_{s-1}]_{\alpha}{}^{\beta}$.
The original definition of subprincipal symbol goes back to J.J. Duistermaat and L. Hörmander [24, Eqn. (5.2.8)]. They defined the subprincipal symbol for operators acting on half-densities.
@@
The definition of subprincipal symbol (3.4) extends further, without change, to operators acting on $m$-columns of scalar fields (sections of the trivial $\mathbb{R}^{m}$- or $\mathbb{C}^{m}$-bundle over $M$).
A notion of subprincipal symbol for operators acting on 1-forms is available in the literature in the special case when the principal symbol is of the form $fI$, where $f$ is a scalar function on $T^{*}M\setminus\{0\}$ and $I$ is the $d\times d$ identity matrix, see, e.g., [31, 35].
@@
$[Q_{\mathrm{sub}}]_{\mu}{}^{\nu}:=\left([q_{s-1}]_{\mu}{}^{\nu}+\frac{i}{2}\dfrac{\partial^{2}[q_{s}]_{\mu}{}^{\nu}}{\partial x^{\gamma}\partial\xi_{\gamma}}\right)+\frac{i}{2}\left(\Gamma^{\alpha}{}_{\gamma\alpha}\dfrac{\partial[q_{s}]_{\mu}{}^{\nu}}{\partial\xi_{\gamma}}\ -\Gamma^{\alpha}{}_{\gamma\mu}\dfrac{\partial[q_{s}]_{\alpha}{}^{\nu}}{\partial\xi_{\gamma}}-\Gamma^{\nu}{}_{\gamma\alpha}\dfrac{\partial[q_{s}]_{\mu}{}^{\alpha}}{\partial\xi_{\gamma}}\right)$ (3.6)
@@
‘Covariance’ in Theorem 3.1 means that one doesn’t get derivatives of Jacobians under change of local coordinates. More precisely, let $x$ and $\widetilde{x}$ be two local coordinate systems, and let $Q_{\mathrm{sub}}$ and $\widetilde{Q}_{\mathrm{sub}}$ be the quantity (3.6) evaluated in coordinates $x$ and $\widetilde{x}$ respectively. Put $J^{\alpha}{}_{\beta}:=\partial\widetilde{x}^{\alpha}/\partial x^{\beta}$, $[J^{-1}]^{\mu}{}_{\nu}:=\partial x^{\mu}/\partial\widetilde{x}^{\nu}$, $\widetilde{\xi}_{\alpha}:=[J^{-1}]^{\beta}{}_{\alpha}\,\xi_{\beta}\,$. Then the statement of the theorem is that
$[\widetilde{Q}_{\mathrm{sub}}]_{\alpha}{}^{\beta}(\widetilde{x},\widetilde{\xi})=[J^{-1}]^{\mu}{}_{\alpha}(\widetilde{x})\ [Q_{\mathrm{sub}}]_{\mu}{}^{\nu}(x(\widetilde{x}),\xi(\widetilde{x},\widetilde{\xi}))\ J^{\beta}{}_{\nu}(x(\widetilde{x}))\ .$ (3.7)
@@
$[q_{s-1}]_{\mu}{}^{\nu}+\frac{i}{2}\nabla_{\gamma}\dfrac{\partial[q_{s}]_{\mu}{}^{\nu}}{\partial\xi_{\gamma}}$
@@
$(z,\xi)\mapsto\left[S^{-1}(\rho^{1/2}SQS^{-1}\rho^{-1/2})_{\mathrm{sub}}S\right]_{\alpha}{}^{\beta}(z,\xi),$ (3.12)
@@
denotes the usual subprincipal symbol [24, Eqn. (5.2.8)] of the operator $Q_{1/2}:=\rho^{1/2}SQS^{-1}\rho^{-1/2}$ acting on $d$-columns of half-densities.
@@
$[\widetilde{q}_{1/2}]_{j}{}^{k}(x,y,\xi):=\left[\frac{\rho(x)}{\rho(y)}\right]^{1/2}\widetilde{e}_{j}{}^{\alpha}(x)\,q_{\alpha}{}^{\beta}(x,\xi)\,\widetilde{e}\,{}^{k}{}_{\beta}(y)\,.$
@@
The terms positively homogeneous in $\xi$ of degree $s$ and $s-1$ in
$[q_{1/2}]_{j}{}^{k}(x,\xi):=\mathcal{S}_{\mathrm{left}}([\widetilde{q}_{1/2}]_{j}{}^{k}(x,y,\xi))$
@@
$[(q_{1/2})_{s}]_{j}{}^{k}(x,\xi)=\widetilde{e}_{j}{}^{\alpha}(x)\,[q_{s}]_{\alpha}{}^{\beta}(x,\xi)\,\widetilde{e}\,{}^{k}{}_{\beta}(x)$ (3.14)
@@
$[(q_{1/2})_{s-1}]_{j}{}^{k}(x,\xi)=\widetilde{e}_{j}{}^{\alpha}\,[q_{s-1}]_{\alpha}{}^{\beta}\widetilde{e}\,{}^{k}{}_{\beta}-i\widetilde{e}_{j}{}^{\alpha}\dfrac{\partial[q_{s}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\dfrac{\partial\widetilde{e}\,{}^{k}{}_{\beta}}{\partial x^{\gamma}}+\frac{i}{2}(\ln\rho)_{x^{\gamma}}\,\widetilde{e}_{j}{}^{\alpha}\dfrac{\partial[q_{s}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\widetilde{e}\,{}^{k}{}_{\beta}\,,$ (3.15)
@@
$[(Q_{1/2})_{\mathrm{sub}}]_{j}{}^{k}(z,\xi)=\left.\left([(q_{1/2})_{s-1}]_{j}{}^{k}(x,\xi)+\frac{i}{2}\dfrac{\partial^{2}[(q_{1/2})_{s}]_{\alpha}{}^{\beta}}{\partial x^{\gamma}\partial\xi_{\gamma}}(x,\xi)\right)\right|_{x=z}\\ =\widetilde{e}_{j}{}^{\alpha}\,\left([(q_{1/2})_{s-1}]_{\alpha}{}^{\beta}+\frac{i}{2}\dfrac{\partial^{2}[(q_{1/2})_{s}]_{\alpha}{}^{\beta}}{\partial x^{\gamma}\partial\xi_{\gamma}}\right)\widetilde{e}\,{}^{k}{}_{\beta}+\frac{i}{2}\dfrac{\partial\widetilde{e}_{j}{}^{\alpha}}{\partial x^{\gamma}}\dfrac{\partial[(q_{1/2})_{s}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\widetilde{e}\,{}^{k}{}_{\beta}\\ -\frac{i}{2}\widetilde{e}_{j}{}^{\alpha}\dfrac{\partial[(q_{1/2})_{s}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\dfrac{\partial\widetilde{e}\,{}^{k}{}_{\beta}}{\partial x^{\gamma}}+\frac{i}{2}(\ln\rho)_{x^{\gamma}}\,\widetilde{e}_{j}{}^{\alpha}\dfrac{\partial[(q_{1/2})_{s}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\widetilde{e}\,{}^{k}{}_{\beta}\,.$ (3.16)
@@
$\left.\widetilde{e}\,{}^{j}{}_{\mu}\dfrac{\partial\widetilde{e}_{j}{}^{\alpha}}{\partial x^{\gamma}}\right|_{x=z}=-\Gamma^{\alpha}{}_{\gamma\mu}(z),\qquad\left.\dfrac{\partial\widetilde{e}\,{}^{k}{}_{\beta}}{\partial x^{\gamma}}\,\widetilde{e}_{k}{}^{\nu}\right|_{x=z}=\Gamma^{\nu}{}_{\gamma\beta}(z),$ (3.17)

### v4

@@ 3 Pseudodifferential operators acting on 1-forms
[…]
@@
[…]
@@
However, the subleading term, $q_{s-1}]_{\alpha}{}^{\beta}\,$, is not invariant under change of coordinates. The task at hand is to define and provide a formula for the subprincipal symbol of $Q$, by determining appropriate correction terms to $q_{s-1}]_{\alpha}{}^{\beta}$.
The original definition of subprincipal symbol goes back to J.J. Duistermaat and L. Hörmander [26, Eqn. (5.2.8)]. They defined the subprincipal symbol for operators acting on half-densities.
@@
The definition of subprincipal symbol (3.4) extends further, without change, to operators acting on $m$-columns of scalar fields (sections of the trivial $\mathbb{R}^{m}$- or $\mathbb{C}^{m}$-bundle over $M$). By choosing a global framing (an orientable $3$-manifold is automatically parallelizable), one can then produce a coordinate-invariant reformulation of the classical construction of subprincipal symbol for operators on 1-forms. However, we should like to emphasise that the approach presented in this manuscript avoids the reliance on a specific global framing. For this reason, we believe it is still of interest and worth presenting.
Note that a notion of subprincipal symbol for operators acting on 1-forms is available in the literature in the special case when the principal symbol is of the form $fI$, where $f$ is a scalar function on $T^{*}M\setminus\{0\}$ and $I$ is the $d\times d$ identity matrix, see, e.g., [36, 40].
@@
$Q_{\mathrm{sub}}]_{\mu}{}^{\nu}:=\left([q_{s-1}]_{\mu}{}^{\nu}+\frac{i}{2}\dfrac{\partial^{2}[q_{s}]_{\mu}{}^{\nu}}{\partial x^{\gamma}\partial\xi_{\gamma}}\right)+\frac{i}{2}\left(\Gamma^{\alpha}{}_{\gamma\alpha}\dfrac{\partial[q_{s}]_{\mu}{}^{\nu}}{\partial\xi_{\gamma}}\ -\Gamma^{\alpha}{}_{\gamma\mu}\dfrac{\partial[q_{s}]_{\alpha}{}^{\nu}}{\partial\xi_{\gamma}}-\Gamma^{\nu}{}_{\gamma\alpha}\dfrac{\partial[q_{s}]_{\mu}{}^{\alpha}}{\partial\xi_{\gamma}}\right)$ (3.6)
@@
‘Covariance’ in Theorem 3.1 means that one doesn’t get derivatives of Jacobians under change of local coordinates. More precisely, let $x$ and $\widetilde{x}$ be two local coordinate systems, and let $Q_{\mathrm{sub}}$ and $\widetilde{Q}_{\mathrm{sub}}$ be the quantity (3.6) evaluated in coordinates $x$ and $\widetilde{x}$ respectively. Put $J^{\alpha}{}_{\beta}:=\partial\widetilde{x}^{\alpha}/\partial x^{\beta}$, $J^{-1}]^{\mu}{}_{\nu}:=\partial x^{\mu}/\partial\widetilde{x}^{\nu}$, $\widetilde{\xi}_{\alpha}:=[J^{-1}]^{\beta}{}_{\alpha}\,\xi_{\beta}\,$. Then the statement of the theorem is that
[…]
@@
$q_{s-1}]_{\mu}{}^{\nu}+\frac{i}{2}\nabla_{\gamma}\dfrac{\partial[q_{s}]_{\mu}{}^{\nu}}{\partial\xi_{\gamma}}$
@@
[…]
@@
denotes the usual subprincipal symbol [26, Eqn. (5.2.8)] of the operator $Q_{1/2}:=\rho^{1/2}SQS^{-1}\rho^{-1/2}$ acting on $d$-columns of half-densities.
@@
[…]
@@
The terms positively homogeneous in $\xi$ of degree $s$ and s1s-1 in
$)[q_{1/2}]_{j}{}^{k}(x,\xi):=\mathcal{S}_{\mathrm{left}}([\widetilde{q}_{1/2}]_{j}{}^{k}(x,y,\xi$))
@@
[…]
@@
$(q_{1/2})_{s-1}]_{j}{}^{k}(x,\xi)=\widetilde{e}_{j}{}^{\alpha}\,[q_{s-1}]_{\alpha}{}^{\beta}\widetilde{e}\,{}^{k}{}_{\beta}-i\widetilde{e}_{j}{}^{\alpha}\dfrac{\partial[q_{s}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\dfrac{\partial\widetilde{e}\,{}^{k}{}_{\beta}}{\partial x^{\gamma}}+\frac{i}{2}(\ln\rho)_{x^{\gamma}}\,\widetilde{e}_{j}{}^{\alpha}\dfrac{\partial[q_{s}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\widetilde{e}\,{}^{k}{}_{\beta}\,$, (3.15)
@@
$(Q_{1/2})_{\mathrm{sub}}]_{j}{}^{k}(z,\xi)=\left.\left([(q_{1/2})_{s-1}]_{j}{}^{k}(x,\xi)+\frac{i}{2}\dfrac{\partial^{2}[(q_{1/2})_{s}]_{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}j}{}^{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}k}}{\partial x^{\gamma}\partial\xi_{\gamma}}(x,\xi)\right)\right|_{x=z}\\ =\widetilde{e}_{j}{}^{\alpha}\,\left([{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{q_{s-1}}}]_{\alpha}{}^{\beta}+\frac{i}{2}\dfrac{\partial^{2}[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{q_{s}}}]_{\alpha}{}^{\beta}}{\partial x^{\gamma}\partial\xi_{\gamma}}\right)\widetilde{e}\,{}^{k}{}_{\beta}+\frac{i}{2}\dfrac{\partial\widetilde{e}_{j}{}^{\alpha}}{\partial x^{\gamma}}\dfrac{\partial[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{q_{s}}}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\widetilde{e}\,{}^{k}{}_{\beta}\\ -\frac{i}{2}\widetilde{e}_{j}{}^{\alpha}\dfrac{\partial[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{q_{s}}}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\dfrac{\partial\widetilde{e}\,{}^{k}{}_{\beta}}{\partial x^{\gamma}}+\frac{i}{2}(\ln\rho)_{x^{\gamma}}\,\widetilde{e}_{j}{}^{\alpha}\dfrac{\partial[{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}{q_{s}}}]_{\alpha}{}^{\beta}}{\partial\xi_{\gamma}}\widetilde{e}\,{}^{k}{}_{\beta}\,$. (3.16)
@@
[…]